The viscous drag F experience by a spherical bob of radius (r) falling through a liquid is proportional to the radius of the bob, coefficient of viscosity, and terminal velocity (v). If the complant of propotionality is 6 pie, find the relationship between F, Radius, Viscosity and velocity


1
Expert's answer
2021-01-26T17:48:45-0500
"F=6\pi \eta^a v^bR^c\\ [F]=[\eta]^a [v]^b[R]^c\\ [MLT^{-2}]=[ML^{-1}T^{-1}]^a [LT^{-1}]^b[L]^c\\a=1,b=1,c=1\\"

"F=6\pi \eta vR"